Assume that impure copper contains only Fe, Au, Ag as impurities. After passing of 140 ampere for 482.5s, the mass of anode decreased by 22.260g and the cathode increased in mass by 22.011g. Calculate the percentage of iron and percentage of copper originally present.

Grade:11

1 Answers

Sunil Kumar FP
9 years ago
THE increase in mass of cathode is due to deposition of Cu and the loss in mass of anode is due to loss of Cu and Fe
and loss of Ag and Au to fall in anode mud
gain in weight at cathode due to copper=22.011g
mole=22.011/63.5=.3466

equivalent of Cu and Fe dissolved at anode=it/96500
=140*482.5/96500=.70
mole of Cu and Fe dissolved at anode=.7/2=.35
mole of Fe=.3500-.3466
=.0034

weight of Fe dissolved at anode=.0034*56=.190g
thus anode weight loss of 22.260g contains 22.011gCu,.19gFe,Au+Ag=.059g
%Cu=22.011/22.26=98.88%
%Fe=.190/22.26=.85%